OpenStudy (anonymous):

Find the inverse g(x) of the function 1.f(x)=x-1

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

@Compassionate

OpenStudy (igreen):

@thegoldenvok Just switch the sign: \(f(x) = x - 1\) Inverse: \(f(x)^{-1} = x + 1\)

OpenStudy (anonymous):

so thats the answer?

OpenStudy (igreen):

Yup.